OpenStudy (anonymous):

If b+c=pi prove that 2(1-sinbsinc) = cos(sqr)b+cos(sqr)c?

OpenStudy (anonymous):

LIke this? if b + c = pi, then:\[2(1-\sin b+\sin c)=\cos( \sqrt b)+\cos (\sqrt c)\]?

OpenStudy (anonymous):

cos square b + cos square c

OpenStudy (anonymous):

or squared?

OpenStudy (anonymous):

oh, ok.

OpenStudy (anonymous):

i fixed one little mistake in the question

OpenStudy (anonymous):

man i completely read the statement wrong >.<

OpenStudy (anonymous):

oh, ok, i thought i was going crazy.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Basically, since b + c = pi, we have that sin b = sin c, and we exploit that. I guess I could have done the proof the other way around using that:\[2(1-\sin b \sin c)=(1-\sin b \sin c)+(1-\sin b \sin c)\]\[=(1-\sin ^2b)+(1-\sin ^2c)=\cos^2 b +\cos^2c\]

OpenStudy (anonymous):

THANK YOU!